# arXiv:2603.01336v1[math.FA] 02 Mar 2026

### 02 Mar 2026 v1

@@ 1. Introduction
The class of uaw-Dunford-Pettis was defined in [5]. This article introduces and explores previously unstudied properties and provide new useful characterizations of uaw-Dunford-Pettis operator. Highlighting its relationship with other classes of operators. Furthermore, Throughout this work we introduce and investigate the weak* version of this operator, that we called the disjoint weak Dunford-Pettis operators. We provide a several results about this new class of operators, and we establish new characterization of Banach Lattices.
@@ 2. Preliminaries
  • Unbounded absolutely weak Dunford–Pettis operator, abbreviated as (uaw-Dunford–Pettis, shortly), if for every norm bounded sequence $(x_{n})$ in $E$ , $x_{n}\overset{uaw}{\longrightarrow}0$ implies $\|T(x_{n})\|\rightarrow 0$ uaw-Dunford-Pettis, see [5].
@@ 3. Main Results
$T$ is a weak Dunford-Pettis operator (because $\ell^{1}$ has the Dunford-Pettis property), and by hypothesis $T$ is uaw-Dunford-Pettis. Now, let $(e_{n})$ be the standard basis of ∎
@@
$(1)\Longrightarrow(2)$ Let $T:E\longrightarrow c_{0}$ be an arbitrary linear operator. We will prove that $T(x_{n})||\longrightarrow 0$, for every disjoint sequence $(x_{n})\subset E$. Assume by way of contradiction that $(T(x_{n}))||$ does not converge to 0 for some disjoint sequence $(x_{n})\in E$. Then, we can suppose that there would exists some $\epsilon>0$ such that $T(x_{n})||>\epsilon$ for all $n\in\mathbb{N}$.
@@
Now, we consider some ideal properties for the class of disjoint weak Dunford-Pettis the operator. But before investigating this, we give the following Example.
@@
$(1)$ Let $(x_{n})$ be a disjoint sequence of $E$ and $(f_{n})\subset G^{\prime}$ such that $f_{n}\overset{w^{\star}}{\longrightarrow}0$. Then $f_{n}\circ T\overset{w^{\star}}{\longrightarrow}0\quad(\star)$.
@@
$(2)$ Let $(x_{n})$ be a disjoint sequence of $E$ and $(f_{n})\subset G^{\prime}$ such that $f_{n}\overset{w^{\star}}{\longrightarrow}0$. As, $S$ is disjointness preserving operator, it follows that $(S(x_{n}))$ is a disjoint sequence. In this step assumption asserts that $f_{n}\left(T(S(x_{n})\right)\longrightarrow 0$. Therefore, $T\circ S$ is disjoint weak Dunford-Pettis.
@@
Let $(x_{n})$ be a disjoint sequence in $E$, by Lemma 2 of [16], $x_{n}\overset{uaw}{\longrightarrow}0$. Then, by Lemma 1 of [16], $|x_{n}|\overset{uaw}{\longrightarrow}0$. Now, using o[[16],Theorem 7] it follows that $x_{n}\overset{w}{\longrightarrow}0$ and $|x_{n}|\overset{w}{\longrightarrow}0$.
@@
$(1)\Longrightarrow(2)$ Let $(f_{n})\subset E^{\prime}$ such that $f_{n}\overset{w^{\star}}{\longrightarrow}0$, and a norm bounded disjoint sequence $(x_{n})\subset E$.
@@
Assume by way of a contradiction that $T$ is not disjoint weak Dunford-Pettis, that is, there exist $(x_{n})$ a bounded disjoint sequence in $E$ and $(f_{n})$ a sequence of $F^{\prime}$ such that $f_{n}\overset{w*}{\longrightarrow}0$ in $F^{\prime}$ with $f_{n}(Tx_{n})\nrightarrow 0$. Therefore there exists a subsequence of $(f_{n}(Tx_{n}))$ (which we can denote by $(f_{n}(Tx_{n}))$) and $\varepsilon>$ such that $|f_{n}(Tx_{n})|>\varepsilon$, for all $n$. In particular, we infer from the inequality $|f_{n}(Tx_{n})|\leq|f_{n}|(T|x_{n}|)$, that $|f_{n}|(T|x_{n}|)|>\varepsilon$. According to [[16],Theorem 7] $|x_{n}|\overset{w}{\longrightarrow}0$, and hence $T(|x_{n}|)\overset{w}{\longrightarrow}0$. Now, an easy inductive argument shows that there exists a subsequence $(z_{n})$ of $(|x_{n}|)$ and subsequence $(g_{n})$ of $(f_{n})$ such that
@@
In the following, we give necessary conditions for which each weak Dunford-Pettis operator is L-weakly compact.
@@
So, there exists a subsequence of $(y_{k})$, that we denote without loss of generality by $(y_{k})$, and $y\in E$ such that $y_{k}-y\overset{w}{\rightarrow}0$. On the other hand, by [10, Corollary 2.4.3], it follows that $g_{n}\overset{w*}{\longrightarrow}0$. By hypothesis $T$ is disjoint weak Dunford-Pettis, hence $|g_{k}(Tz_{k})|\rightarrow 0$. Now, as $F$ has the DP* property it follows that $|g_{k}(T(y_{k}-y))|\rightarrow 0$, and we have also $|g_{k}(T(y))|\rightarrow 0$, as $k\rightarrow\infty$, combining all and from the inequality
@@
$1\Longrightarrow 2$ Let $A$ be a $\sigma$-uaw-compact subset of $E$, we need to show that $T(A)$ is limited subset of $F$. To this end, let $f_{n}\overset{w*}{\longrightarrow}0$ in $F^{\prime}$, let $y_{n}\in T(A)$, that is, there exists $x_{n}\in A$ such that $y_{n}=T(x_{n})$. Since $A$ is $\sigma$-uaw compact, there exists $(x_{n_{k}})$ a subsequence of $(x_{n})$ and $x\in E$ such that $x_{n_{k}}\overset{uaw}{\longrightarrow}x$, we may assume that $x=0$. By Theorem 8.11 [14], there exists $(x_{n_{k^{\prime}}})$ a subsequence of $(x_{n_{k}})$ a subsequence and a disjoint sequence $(d_{k^{\prime}})$ satisfying $x_{n_{k^{\prime}}}-d_{k^{\prime}}\overset{w}{\longrightarrow}0$, therefore $T(x_{n_{k^{\prime}}}-d_{k^{\prime}})\overset{w}{\longrightarrow}0$. From one hand, as $F$ has DP* property, $f_{k^{\prime}}(T(x_{n_{k^{\prime}}}-d_{k^{\prime}}))\longrightarrow 0$. On the other hand, by hypothesis $T$ is disjoint weak Dunford-Pettis, that yields $f_{k^{\prime}}(Td_{k^{\prime}})\longrightarrow 0$. Hence $f_{k^{\prime}}(Tx_{n_{k^{\prime}}})\longrightarrow 0$. By the same argument, we showed that every subsequence of $(f_{n}(Tx_{n}))$ has a further subsequence which converges to zero, therefore $f_{n}(Tx_{n})\longrightarrow 0$, which proves that $T(A)$ is limited subset .
@@
Let $E$ and $F$ be a pair of Banach lattices, such that $E$ order continuous. The following are equivalent:
@@
$2\Longrightarrow 1$ Let $(x_{n})$ be a disjoint sequence of $E$, it follows from [[16], Lemma 2 and Theorem 4] that $W=\{{x_{n},n\in\mathbb{N}\}}\cup\{0\}$, is $\sigma$-un-compact subset of $E$, by hypothesis $T(W)$ is limited in $F$. Therefore, for an arbitrary sequence $(f_{n})$ in $F^{\prime}$ such that $f_{n}\overset{w*}{\longrightarrow}0$, we obtain $f_{n}(Tx_{n})\longrightarrow 0$, which shows that $T$ is disjoint weak Dunford-Pettis.

### 19 Apr 2026 v2

# Some properties of uaw-Dunford-Pettis operators

@@ 1. Introduction
The class of uaw-Dunford-Pettis was defined in [5]. This article introduces and explores previously unstudied properties and provide new useful characterizations of uaw-Dunford-Pettis operator. Highlighting its relationship with other classes of operators. Furthermore, Throughout this work we introduce and investigate the weak* version of this operator, that we called the disjoint weak Dunford-Pettis operators. We provide several results about this new class of operators, and we establish new characterization of Banach Lattices.
@@ 2. Preliminaries
  • Unbounded absolutely weak Dunford-Pettis operator, (uaw-Dunford-Pettis, shortly), if for every norm bounded sequence $(x_{n})$ in $E$, $x_{n}\overset{uaw}{\longrightarrow}0$ implies $\|T(x_{n})\|\rightarrow 0$, see [5].
@@ 3. Main Results
$T$ is a weak Dunford-Pettis operator (because $\ell^{1}$ has the Dunford-Pettis property), and by hypothesis $T$ is uaw-Dunford-Pettis. Now, let $(e_{n})$ be the standard basis of $\ell^{1}$, we have $e_{n}\overset{uaw}{\longrightarrow}0$, then $\|T(e_{n})\|=\|y_{n}\|\rightarrow 0$. Therefore, $Y$ has the Schur property. ∎
@@
$(1)\Longrightarrow(2)$ Let $T:E\longrightarrow c_{0}$ be an arbitrary linear operator. We will prove that $T(x_{n})||\longrightarrow 0$, for every norm bounded disjoint sequence $(x_{n})\subset E$. Assume by way of contradiction that $(T(x_{n}))||$ does not converge to 0 for some disjoint sequence $(x_{n})\in E$. Then, we can suppose that there would exists some $\epsilon>0$ such that $T(x_{n})||>\epsilon$ for all $n\in\mathbb{N}$.
@@
Now, we consider some ideal properties for the class of disjoint weak Dunford-Pettis operator. But before investigating this, we give the following Example.
@@
$(1)$ Let $(x_{n})$ be a norm bounded disjoint sequence of $E$ and $(f_{n})\subset G^{\prime}$ such that $f_{n}\overset{w^{\star}}{\longrightarrow}0$. Then $f_{n}\circ T\overset{w^{\star}}{\longrightarrow}0\quad(\star)$.
@@
$(2)$ Let $(x_{n})$ be a disjoint sequence of $E$ and $(f_{n})\subset G^{\prime}$ such that $f_{n}\overset{w^{\star}}{\longrightarrow}0$. As, $S$ is disjointness preserving operator, it follows that $(S(x_{n}))$ is a disjoint sequence. In this step, assumption asserts that $f_{n}\left(T(S(x_{n})\right)\longrightarrow 0$. Therefore, $T\circ S$ is disjoint weak Dunford-Pettis.
@@
Let $(x_{n})$ be a disjoint sequence in $E$, by Lemma 2 of [16], $x_{n}\overset{uaw}{\longrightarrow}0$. Then, by Lemma 1 of [16], $|x_{n}|\overset{uaw}{\longrightarrow}0$. Now, using [[16],Theorem 7] it follows that $x_{n}\overset{w}{\longrightarrow}0$ and $|x_{n}|\overset{w}{\longrightarrow}0$.
@@
$(1)\Longrightarrow(2)$ Let $(f_{n})\subset E^{\prime}$ such that $f_{n}\overset{w^{\star}}{\longrightarrow}0$, and let consider a norm bounded disjoint sequence $(x_{n})\subset E$.
@@
Assume by way of a contradiction that $T$ is not disjoint weak Dunford-Pettis, that is, there exist $(x_{n})$ a bounded disjoint sequence in $E$ and $(f_{n})$ a sequence of $F^{\prime}$ such that $f_{n}\overset{w^{\star}}{\longrightarrow}0$ in $F^{\prime}$ with $f_{n}(Tx_{n})\nrightarrow 0$. Therefore there exists a subsequence of $(f_{n}(Tx_{n}))$ (which we can denote by $(f_{n}(Tx_{n}))$) and $\varepsilon>$ such that $|f_{n}(Tx_{n})|>\varepsilon$, for all $n$. In particular, we infer from the inequality $|f_{n}(Tx_{n})|\leq|f_{n}|(T|x_{n}|)$, that $|f_{n}|(T|x_{n}|)|>\varepsilon$. According to [[16],Theorem 7] $|x_{n}|\overset{w}{\longrightarrow}0$, and hence $T(|x_{n}|)\overset{w}{\longrightarrow}0$. Now, an easy inductive argument shows that there exists a subsequence $(z_{n})$ of $(|x_{n}|)$ and subsequence $(g_{n})$ of $(f_{n})$ such that
@@
In the following, we give necessary conditions for which each disjoint weak Dunford-Pettis operator is L-weakly compact.
@@
So, there exists a subsequence of $(y_{k})$, that we denote without loss of generality by $(y_{k})$, and $y\in E$ such that $y_{k}-y\overset{w}{\rightarrow}0$. On the other hand, by [10, Corollary 2.4.3], it follows that $g_{n}\overset{w^{\star}}{\longrightarrow}0$. By hypothesis $T$ is disjoint weak Dunford-Pettis, hence $|g_{k}(Tz_{k})|\rightarrow 0$. Now, as $F$ has the DP* property it follows that $|g_{k}(T(y_{k}-y))|\rightarrow 0$, and we have also $|g_{k}(T(y))|\rightarrow 0$, as $k\rightarrow\infty$, combining all and from the inequality
@@
$1\Longrightarrow 2$ Let $A$ be a $\sigma$-uaw-compact subset of $E$, we need to show that $T(A)$ is limited subset of $F$. To this end, let $f_{n}\overset{w^{\star}}{\longrightarrow}0$ in $F^{\prime}$, let $y_{n}\in T(A)$, that is, there exists $x_{n}\in A$ such that $y_{n}=T(x_{n})$. Since $A$ is $\sigma$-uaw-compact, there exists $(x_{n_{k}})$ a subsequence of $(x_{n})$ and $x\in E$ such that $x_{n_{k}}\overset{uaw}{\longrightarrow}x$, we may assume that $x=0$. By Theorem 8.11 [14], there exists $(x_{n_{k^{\prime}}})$ a subsequence of $(x_{n_{k}})$ a subsequence and a disjoint sequence $(d_{k^{\prime}})$ satisfying $x_{n_{k^{\prime}}}-d_{k^{\prime}}\overset{w}{\longrightarrow}0$, therefore $T(x_{n_{k^{\prime}}}-d_{k^{\prime}})\overset{w}{\longrightarrow}0$. From one hand, as $F$ has DP* property, $f_{k^{\prime}}(T(x_{n_{k^{\prime}}}-d_{k^{\prime}}))\longrightarrow 0$. On the other hand, by hypothesis $T$ is disjoint weak Dunford-Pettis, that yields $f_{k^{\prime}}(Td_{k^{\prime}})\longrightarrow 0$. Hence $f_{k^{\prime}}(Tx_{n_{k^{\prime}}})\longrightarrow 0$. By the same argument, we showed that every subsequence of $(f_{n}(Tx_{n}))$ has a further subsequence which converges to zero, therefore $f_{n}(Tx_{n})\longrightarrow 0$, which proves that $T(A)$ is limited subset .
@@
Let $E$ and $F$ be a pair of Banach lattices, such that $E$ is order continuous. The following are equivalent:
@@
$2\Longrightarrow 1$ Let $(x_{n})$ be a disjoint sequence of $E$, it follows from [[16], Lemma 2 and Theorem 4] that $W=\{{x_{n},n\in\mathbb{N}\}}\cup\{0\}$, is $\sigma$-un-compact subset of $E$, by hypothesis $T(W)$ is limited in $F$. Therefore, for an arbitrary sequence $(f_{n})$ in $F^{\prime}$ such that $f_{n}\overset{w^{\star}}{\longrightarrow}0$, we obtain $f_{n}(Tx_{n})\longrightarrow 0$, which shows that $T$ is disjoint weak Dunford-Pettis.